# Nauplius Optimisation for Autonomous Hydrodynamics

### v1

@@ Abstract
Autonomous Underwater vehicles must operate in strong currents, limited acoustic bandwidth, and persistent sensing requirements where conventional swarm optimisation methods are unreliable. This paper presents Nauplius Optimisation for Autonomous Hydrodynamics (NOAH), a novel nature-inspired swarm optimisation algorithm that combines current-aware drift, irreversible settlement in persistent sensing nodes, and colony-based communication. Drawing inspiration from the behaviour of barnacle nauplii, NOAH addresses the critical limitations of existing swarm algorithms by providing hydrodynamic awareness, irreversible anchoring mechanisms, and colony-based communication capabilities essential for underwater exploration missions. The algorithm establishes a comprehensive foundation for scalable and energy-efficient underwater swarm robotics with validated performance analysis. Validation studies demonstrate an 86% success rate for permanent anchoring scenarios, providing a unified formulation for hydrodynamic constraints and irreversible settlement behaviours with an empirical study under flow.
@@ I Introduction
Underwater exploration and environmental monitoring are increasingly crucial in marine conservation, offshore resource management, and scientific research. The aquatic domain presents formidable challenges due to complex hydrodynamics, limited communication, and severe energy limits [1].
Swarm robotics with Autonomous Underwater Vehicle (AUV)s offers improved coverage, fault tolerance, and adaptability for such tasks [2]. Multiple AUVs operating collectively can survey larger areas faster and withstand single-node failures better than a single vehicle [3]. Recent advances in multi-AUV exploration underscore these benefits, with new distributed strategies achieving robust, wide-area ocean surveys [4].
Wang et al. demonstrate autonomous cooperative systems for underwater target detection and localisation, highlighting the potential for coordinated multi-AUV operations [5]. Wen et al. address intelligent decision-making for the planning of the AUV path against ocean current disturbances [6].
However, swarm optimisation algorithms, originally designed for terrestrial or aerial robots, often fail in aquatic settings due to three critical limitations: difficulty modelling dynamic ocean currents, lack of sustained sensing mechanisms for long-term observation, and challenges with acoustic communication bandwidth and latency constraints [2]. Standard optimisers such as Particle Swarm Optimisation (PSO) and Ant Colony Optimisation (ACO) assume fully mobile agents and simple communication [7], neglecting the fluid forces and intermittent acoustic links that AUV swarms face [8]. Moreover, existing multi-AUV systems rarely allow vehicles to become stationary observers, with irreversible anchoring largely absent except in heterogeneous specialised networks [8]. Although emerging research addresses individual aspects such as energy-efficient AUV deployment [9] and ocean flow-based coverage planning [4], no previous work has merged hydrodynamic drift adaptation, persistent anchoring, and colony-style communication into a unified framework.
@@ II Related Work
PSO Flocking (birds/fish) Velocity + best memory [10], [11]
ACO Ant foraging Pheromone-guided path search [12], [13]
Artificial Bee Colony (ABC) Bee foraging Employed/onlooker/scout phases [14], [15]
Differential Evolution (DE) Evolutionary vectors Mutation & recombination [16]
Firefly Algorithm (FA) Firefly attraction Distance-based movement [17]
Hippopotamus Optimisation Algorithm (HOA) Hippopotamus behaviour Trinary-phase search [18]
Tetragonula Bee Hive construction Distributed coordination [19]
Modified Particle Whale Optimisation Algorithm (MP-WOA) Whale predation Encircling & spiral [20]
Selective Opposition-Based Barnacle Mating Optimisation (SOBMO) Barnacle mating Opposition-based search [21]
Adaptive Bacterial Colony Optimisation (ABCO) Bacterial foraging Explore-exploit-reproduce [22]
@@ II-A Foundational Swarm Optimisation Methods
Foundational swarm algorithms have established the core principles of collective intelligence, but exhibit notable limitations for aquatic environments. Particle Swarm Optimisation (PSO[10] simulates flocking behaviour by updating positions based on the best individual and global positions, finding applications in AUV path planning [11]. Recent work by Li et al. demonstrates improved PSO for AUV 3D path planning [23], while Sun et al. address energy-optimised path planning with ocean current considerations [24]. Ant Colony Optimisation (ACO[12] uses pheromone-mediated foraging for efficient energy routing. Recent advances in dual-strategy ACO algorithms, originally developed for Unmanned Aerial Vehicle (UAV) path planning [13], demonstrate improved convergence and 3D navigation capabilities that can be adapted for AUV applications in underwater environments, with Yuan et al. integrating ACO with dynamic window algorithms for communication efficiency [25]. Ronghua et al. present an improved ACO for the safe planning of AUV paths [26]. Artificial Bee Colony (ABC[14] divides agents into employed, onlooker, and scout bees for the AUV way point traversal [15], with Prasath et al. applying Enhanced ABC algorithms for underwater image enhancement [27]. Differential Evolution (DE[16] uses mutation and recombination for AUV rendezvous planning [11], while Chen et al. address finite-time velocity-free rendezvous control [28]. The Firefly Algorithm (FA[17] models the attraction based on brightness to plan underwater paths [11].
Despite their effectiveness in terrestrial applications, these foundational swarm intelligence methods have been adapted for underwater exploration [29], but lack essential marine-specific capabilities: hydrodynamic flow field adaptation, irreversible settlement mechanisms, colony-based communication protocols and energy-efficient anchoring strategies.
@@ II-B Contemporary Bio-Inspired Approaches
Recent biologically-motivated algorithms introduce novel behavioural metaphors but share fundamental limitations with foundational methods. The Hippopotamus Optimisation Algorithm (HOA[18] models the positioning, defence, and evasion phases, while the Tetragonula Bee Hive Optimisation [19] emphasises distributed coordination through hive construction behaviours. Enhanced whale optimisation [20] strengthens encircling and spiral search patterns, with Yan et al. applying whale optimisation for AUV path planning [30] and Nadimi-Shahraki et al. providing systematic reviews of whale optimisation algorithms [31]. Selective opposition-based barnacle mating optimisation (SOBMO[21] employs barnacle-inspired mating strategies, while Darvishpoor et al. review nature-inspired algorithms from oceans to space [32]. Adaptive bacterial colony optimisation (ABCO[22] combines exploration, exploitation, and reproduction phases for accelerated convergence, and Kaveripakam et al. applied clustering-based dragonfly optimisation for underwater wireless sensor networks [33].
@@ II-C Recent Advances in Underwater Swarm Robotics
Recent developments in underwater swarm robotics have focused on addressing marine environment challenges, although gaps remain in persistent monitoring and environmental adaptation. Connor et al. [2] provide a comprehensive review highlighting critical limitations in acoustic communication bandwidth and latency constraints. Cai et al. [3] introduce multi-AUV distributed collaborative search methods, while Lončar et al. [8] present heterogeneous robotic swarms incorporating static nodes alongside mobile AUVs. Xu et al. [9] address the energy-efficient 3D deployment of AUV-enabled mobile relays, and Mu and Gao [4] develop multi-AUV coverage planners that account for ocean flow and sonar constraints.
@@ III Background
To address these limitations, we propose NOAH (Nauplius Optimisation for Autonomous Hydrodynamics), a novel nature-inspired swarm algorithm designed specifically for marine robotics applications. The algorithm draws inspiration from barnacle nauplii that drift with currents, settle irrevocably in optimal hosts, and form communicating colonies [34]. By fusing these biologically-motivated behaviours, NOAH enables a swarm of AUVs to dynamically adapt to ocean currents, anchor certain members as persistent monitoring nodes, and maintain a robust networked colony.
@@
Problem Formulation Unlike classical PSO/ACO which aims to continuously move agents toward optima, our goal is to convert mobile swarm agents into stationary sensor nodes under hydrodynamic drift constraints. This transforms the problem from continuous mobility optimisation to irreversible deployment under ocean currents.
@@ IV-A Understanding Nauplius Behaviour
Barnacle nauplii undergo a critical developmental transition from free-swimming larvae to cyprid larvae that exhibit irreversible settlement behaviour [34]. This irreversible anchoring mechanism is essential for survival, as cyprids secrete permanent cement that permanently bonds them to suitable substrates [35]. The irreversible nature of settlement ensures that once a cyprid commits to a location, it cannot relocate, making settlement decisions crucial for long-term survival and reproductive success. This biological constraint provides fundamental insights for designing underwater robotic systems that must balance exploration with permanent deployment strategies in dynamic marine environments where relocation is energetically costly and communication is intermittent [36].
@@ IV-B Larval Drift and Hydrodynamic Cues
Barnacle larvae progress through six naupliar stages before reaching the cyprid phase, consistent with observations in both intertidal and epizoic barnacle species. The development time is temperature-dependent, with cyprids emerging in approximately 7 to 8 days at 20 to 25 °C, followed by settlement and metamorphosis reaching juvenile form on day 12 to 18. During this phase of planktonic drift, larvae are passively transported by ambient currents [34], enabling broad dispersal before settlement.
When approaching large marine hosts, hydrodynamics strongly shapes micro-habitats of encounter and settlement. The larvae actively swim to exploit the favourable flow structures for settlement [37]. Recent work by Reidenbach et al. demonstrates hydrodynamic interactions with coral topography and their impact on larval settlement [38], while Ella and Genin examine capture dynamics at different flow speeds [39].
Analyses on small cetaceans show that the commensal barnacle Xenobalanus globicipitis concentrates along the trailing edges of the appendages, with preference for the centre of the fluke. In delphinids, the densities of the dorsal side exceed the densities of the ventral side, consistent with asymmetric fluke kinematics and adhesion-related shear forces [40, 41]. The trailing-edge colonisation pattern can be observed directly on the dorsal fins of killer whales, as shown in Fig. 4.
Figure 4: Photographs of four killer whales (O. orca) observed off northern Baffin Island in the eastern Canadian Arctic with Xenobalanus barnacles attached to the trailing edge of their upper dorsal fins. Adapted from Matthews [42].
@@ IV-C Host Selection and Settlement
Beyond hydrodynamic guidance, chemical confirmation is a key determinant of settlement. Recent work identified a waterborne settlement pheromone for barnacles as adenosine. Settlement activity remains robust to experimental warming and acidification, underscoring the reliability of chemical signalling under fluctuating conditions [43].
Recent studies demonstrate that exogenous adenosine significantly increases cyprid settlement rates and exploration frequency. It also improves the mechanical properties of adhesive footprints [44]. Liu et al. examine chemical signalling in biofilm-mediated bio-fouling [45], providing information on settlement mechanisms.
In practice, cyprids integrate hydrodynamic concentration near suitable micro-habitats with host-derived chemical signals [36]. Settlement occurs only when cue strength and contact time exceed a threshold, as anchoring is irreversible.
@@ IV-D Colony Formation and Influence
Following anchoring, local hydrodynamics and conspecific signalling reinforce gregarious settlement. This drives colony growth along flow-favourable edges [46]. Positive feedback improves reproductive opportunity and feeding efficiency through consistent flow delivery. It imposes negligible drag costs for large hosts compared to body mass.
Host specificity and adaptive evolution in settlement behaviour further enhance colony formation, as demonstrated in coral-associated barnacle larvae [47]. Colonisation patterns such as those shown in Fig. 4 provide the biological basis for NOAH’s anchoring and colony-communication modules.
@@ V-A State and Fields
which combines energy cost, risk exposure, coverage gaps, and information gain into a single score [48].
@@
where $\tilde{v}_{i,t+1}$ is the tentative velocity update. The prefactor $(1-a_{i,t})$ ensures settled agents remain frozen, implementing irreversible anchoring analogous to cyprid larvae [49]. This progressive freeze strategy transitions agents from exploration to exploitation [50, 51], extending multi-swarm PSO frameworks [52] with irreversible anchoring.
@@ V-B Drift and Update Equations
Anchored colonies emit a spatial field that is repulsive in the short range (to avoid overcrowding) [53], but attractive in the longer range (to draw agents into the vicinity). This bio-fouling pressure field mimics how barnacle colonies signal both attraction and exclusion zones, similar to how real barnacles use chemical cues to guide larval settlement [35]. The field incorporates communication noise through probabilistic signal transmission (Equation 4), defined as
@@
where $A\in[0.1,1.0]$ and $B\in[0.1,0.5]$ control the attraction and repulsion strengths, while $\sigma_{a}\in[0.2,0.8]$ and $\sigma_{r}\in[0.05,0.3]$ define the spatial scales (with $\sigma_{a}>\sigma_{r}$). The communication parameters $R_{k}\in[0.1,0.5]$ represent the reliable communication range for the colony $k$, and $\alpha\in[1.0,5.0]$ controls signal attenuation beyond this range, reflecting realistic underwater acoustic propagation characteristics [2].
The parameter ranges reflect biological observations. Attraction strength $A$ corresponds to chemical signalling intensity in barnacle colonies, while repulsion strength $B$ models physical crowding constraints. The spatial scales $\sigma_{a}$ and $\sigma_{r}$ mirror the effective range of chemical signals versus the zones of immediate physical contact observed in marine larvae [36].
Based on cyprid swimming speeds of $\sim$1.8 cm/s [37, 35], the attraction ranges significantly exceed the immediate contact zones. This enables larvae to detect and navigate the appropriate settlement sites from considerable distances. The dual-scale approach ensures that agents are drawn toward productive areas, but avoid overcrowding. This mirrors the biological phenomenon where barnacle larvae are attracted to established colonies but settle at appropriate distances to avoid competition for resources.
The gradient $G_{\Phi}(x)=\nabla\Phi(x)$ acts as a virtual force on free agents, guiding their movement toward optimal settlement regions while maintaining spatial diversity [54, 55]. This dual-scale interaction mirrors empirical observations in barnacle larvae, where cyprids are chemically attracted to productive conspecific aggregations, but still reject immediate-contact regions to avoid overcrowding [56]. Furthermore, larvae can detect dissolved settlement signals under flow before making direct contact with the substrate, allowing field-driven navigation rather than purely contact-based adhesion [57]. The Gaussian-based potential field approach follows established methods of the artificial potential field for swarm robotics [58], while the communication probability model incorporates realistic characteristics of the wireless link observed in transition regions [59]. The attraction-repulsion dynamics are based on social force models [60] (see Equation 4).
@@
The velocity update mechanism integrates five biologically inspired components that mirror how barnacle nauplii navigate within hydrodynamic environments. Let $\hat{\nabla}f(x_{i,t})$ denote a surrogate estimate of the local gradient of $f$ at the position of agent $i$. This estimate may be obtained through finite differences or relative fitness ranks while keeping NOAH derivative-free. The tentative velocity update $\tilde{v}_{i,t+1}$ is given by:
@@
$\quad+\underbrace{\beta\left(-\hat{\nabla}f(x_{i,t})\right)}_{\text{fitness gradient}}+\underbrace{\gamma U(x_{i,t})}_{\text{drift (flow)}}$
@@
where $\xi_{i,t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ introduces stochastic exploration. The inertia term $(\omega v_{i,t})$ follows the particle swarm optimisation framework [61], where the inertia weight $\omega$ balances global and local search capabilities [62]. The random component $(\eta\xi_{i,t})$ employs Gaussian perturbations consistent with stochastic approximation theory, particularly simultaneous perturbation stochastic approximation methods [63, 64]. The fitness gradient term $(\beta(-\hat{\nabla}f))$ uses finite difference approximations for derivative-free optimisation [65, 66], while the drift term $(\gamma U(x_{i,t}))$ models the hydrodynamic flow effects observed in ocean sampling applications [67]. Finally, the colony field $(\delta G_{\Phi}(x_{i,t}))$ incorporates the mean-field control principles for swarm coordination [68, 69].
The parameter ranges are biologically motivated based on observed swimming speeds of 2.6 to 15.2 body lengths per second [70] and rheotactic attachment behaviours in which 93% of cyprids swim upstream before settlement [37]. Momentum retention $\omega\in[0.5,0.9]$ reflects larval inertia, exploration parameter $\eta\in[0.1,0.5]$ models stochastic swimming variability, gradient following strength $\beta\in[0.2,1.0]$ corresponds to chemotactic responsiveness, current adaptation $\gamma\in[0.3,0.8]$ reflects hydrodynamic flow exploitation, and colony influence $\delta\in[0.1,0.6]$ balances attraction to settlement-rich regions with localised spacing effects [40].

### v2

@@ Abstract
Autonomous Underwater vehicles must operate in strong currents, limited acoustic bandwidth, and persistent sensing requirements where conventional swarm optimisation methods are unreliable. This paper formulates an irreversible hydrodynamic deployment problem for Autonomous Underwater Vehicle (AUV) swarms and presents Nauplius Optimisation for Autonomous Hydrodynamics (NOAH), a novel nature-inspired swarm optimisation algorithm that combines current-aware drift, irreversible settlement in persistent sensing nodes, and colony-based communication. Drawing inspiration from the behaviour of barnacle nauplii, NOAH addresses the critical limitations of existing swarm algorithms by providing hydrodynamic awareness, irreversible anchoring mechanisms, and colony-based communication capabilities essential for underwater exploration missions. The algorithm establishes a comprehensive foundation for scalable and energy-efficient underwater swarm robotics with validated performance analysis. Validation studies demonstrate an 86% success rate for permanent anchoring scenarios, providing a unified formulation for hydrodynamic constraints and irreversible settlement behaviours with an empirical study under flow.
@@ I Introduction
Underwater exploration and environmental monitoring are increasingly crucial in marine conservation, offshore resource management and scientific research. The aquatic domain presents formidable challenges due to complex hydrodynamics, limited communication and severe energy limitations [Hasan2024].
Swarm robotics with AUVs offers improved coverage, fault tolerance, and adaptability for such tasks [connor2021current]. Multiple AUVs operating collectively can survey larger areas faster and withstand single-node failures better than a single vehicle [cai2024multi]. Recent advances in multi-AUV exploration underscore these benefits, with new distributed strategies achieving robust, wide-area ocean surveys [mu2025coverage].
Wang et al. demonstrate autonomous cooperative systems for underwater target detection and localisation, highlighting the potential for coordinated multi-AUV operations [wang2023multi]. Wen et al. address intelligent decision-making for the planning of the AUV path against ocean current disturbances [wen2024intelligent].
However, swarm optimisation algorithms, originally designed for terrestrial or aerial robots, often fail in aquatic settings due to three critical limitations: difficulty modelling dynamic ocean currents, lack of sustained sensing mechanisms for long-term observation, and challenges with acoustic communication bandwidth and latency constraints [connor2021current]. Standard optimisers such as Particle Swarm Optimisation (PSO) and Ant Colony Optimisation (ACO) assume fully mobile agents and simple communication [Lhotska2006], neglecting the fluid forces and intermittent acoustic links that AUV swarms face [loncar2019heterogeneous]. Moreover, existing multi-AUV systems rarely allow vehicles to become stationary observers, with irreversible anchoring largely absent except in heterogeneous specialised networks [loncar2019heterogeneous]. Although emerging research addresses individual aspects such as energy-efficient AUV deployment [xu2025energy] and ocean flow-based coverage planning [mu2025coverage], no previous work has merged hydrodynamic drift adaptation, persistent anchoring, and colony-style communication into a unified framework. In this context, we address irreversible hydrodynamic deployment by converting mobile AUV swarms into permanently anchored sensing and relay colonies under ocean flow, and we validate this new NOAH framework through anchoring accuracy, convergence, and comparative performance experiments.
@@ II Related Work
PSO Flocking (birds/fish) Velocity + best memory [kennedy1995particle], [panda2020comprehensive]
ACO Ant foraging Pheromone-guided path search [dorigo2006ant], [mai2023uav]
Artificial Bee Colony (ABC) Bee foraging Employed/onlooker/scout phases [karaboga2005idea], [contreras2015mobile]
Differential Evolution (DE) Evolutionary vectors Mutation & recombination [storn1997differential]
Firefly Algorithm (FA) Firefly attraction Distance-based movement [yang2010firefly]
Hippopotamus Optimisation Algorithm (HOA) Hippopotamus behaviour Trinary-phase search [amiri2024hippopotamus]
Tetragonula Bee Hive construction Distributed coordination [gamez2025novel]
Modified Particle Whale Optimisation Algorithm (MP-WOA) Whale predation Encircling & spiral [chakraborty2021novel]
Selective Opposition-Based Barnacle Mating Optimisation (SOBMO) Barnacle mating Opposition-based search [ahmed2024selective]
Adaptive Bacterial Colony Optimisation (ABCO) Bacterial foraging Explore-exploit-reproduce [kogam2025adaptive]
@@ II-A Foundational Swarm Optimisation Methods
Foundational swarm algorithms have established the core principles of collective intelligence, but exhibit notable limitations for aquatic environments. Particle Swarm Optimisation (PSO[kennedy1995particle] simulates flocking behaviour by updating positions based on the best individual and global positions, finding applications in AUV path planning [panda2020comprehensive]. Recent work by Li et al. demonstrates improved PSO for AUV 3D path planning [li2025auv], while Sun et al. address energy-optimised path planning with ocean current considerations [sun2022energy]. Ant Colony Optimisation (ACO[dorigo2006ant] uses pheromone-mediated foraging for efficient energy routing. Recent advances in dual-strategy ACO algorithms, originally developed for Unmanned Aerial Vehicle (UAV) path planning [mai2023uav], demonstrate improved convergence and 3D navigation capabilities that can be adapted for AUV applications in underwater environments, with Yuan et al. integrating ACO with dynamic window algorithms for communication efficiency [yuan2025integrated]. Ronghua et al. present an improved ACO for the safe planning of AUV paths [ronghua2024improved]. Artificial Bee Colony (ABC[karaboga2005idea] divides agents into employed, onlooker, and scout bees for the AUV way point traversal [contreras2015mobile], with Prasath et al. applying Enhanced ABC algorithms for underwater image enhancement [Prasath2020]. Differential Evolution (DE[storn1997differential] uses mutation and recombination for AUV rendezvous planning [panda2020comprehensive], while Chen et al. address finite-time velocity-free rendezvous control [chen2022finite]. The Firefly Algorithm (FA[yang2010firefly] models the attraction based on brightness to plan underwater paths [panda2020comprehensive].
Despite their effectiveness in terrestrial applications, these foundational swarm intelligence methods have been adapted for underwater exploration [Aghababa2012], but lack essential marine-specific capabilities: hydrodynamic flow field adaptation, irreversible settlement mechanisms, colony-based communication protocols and energy-efficient anchoring strategies.
@@ II-B Contemporary Bio-Inspired Approaches
Recent biologically-motivated algorithms introduce novel behavioural metaphors but share fundamental limitations with foundational methods. The Hippopotamus Optimisation Algorithm (HOA[amiri2024hippopotamus] models the positioning, defence, and evasion phases, while the Tetragonula Bee Hive Optimisation [gamez2025novel] emphasises distributed coordination through hive construction behaviours. Enhanced whale optimisation [chakraborty2021novel] strengthens encircling and spiral search patterns, with Yan et al. applying whale optimisation for AUV path planning [yan2021two] and Nadimi-Shahraki et al. providing systematic reviews of whale optimisation algorithms [nadimi2023systematic]. Selective opposition-based barnacle mating optimisation (SOBMO[ahmed2024selective] employs barnacle-inspired mating strategies, while Darvishpoor et al. review nature-inspired algorithms from oceans to space [darvishpoor2023nature]. Adaptive bacterial colony optimisation (ABCO[kogam2025adaptive] combines exploration, exploitation, and reproduction phases for accelerated convergence, and Kaveripakam et al. applied clustering-based dragonfly optimisation for underwater wireless sensor networks [kaveripakam2023clustering].
@@ II-C Recent Advances in Underwater Swarm Robotics
Recent developments in underwater swarm robotics have focused on addressing marine environment challenges, although gaps remain in persistent monitoring and environmental adaptation. Connor et al. [connor2021current] provide a comprehensive review highlighting critical limitations in acoustic communication bandwidth and latency constraints. Cai et al. [cai2024multi] introduce multi-AUV distributed collaborative search methods, while Lončar et al. [loncar2019heterogeneous] present heterogeneous robotic swarms incorporating static nodes alongside mobile AUVs. Xu et al. [xu2025energy] address the energy-efficient 3D deployment of AUV-enabled mobile relays, and Mu and Gao [mu2025coverage] develop multi-AUV coverage planners that account for ocean flow and sonar constraints.
@@ III Background
To address these limitations, we propose NOAH (Nauplius Optimisation for Autonomous Hydrodynamics), a novel nature-inspired swarm algorithm designed specifically for marine robotics applications. The algorithm draws inspiration from barnacle nauplii that drift with currents, settle irrevocably in optimal hosts, and form communicating colonies [nogata2006larval]. By fusing these biologically-motivated behaviours, NOAH enables a swarm of AUVs to dynamically adapt to ocean currents, anchor certain members as persistent monitoring nodes, and maintain a robust networked colony.
@@
Problem Formulation Unlike classical PSO/ACO which aim to continuously move agents toward optima, our goal is to solve an irreversible hydrodynamic deployment problem by converting mobile swarm agents into stationary sensor nodes under hydrodynamic drift constraints. This transforms the problem from continuous mobility optimisation to irreversible deployment under ocean currents.
@@ IV-A Understanding Nauplius Behaviour
Barnacle nauplii undergo a critical developmental transition from free-swimming larvae to cyprid larvae that exhibit irreversible settlement behaviour [nogata2006larval]. This irreversible anchoring mechanism is essential for survival, as cyprids secrete permanent cement that permanently bonds them to suitable substrates [Kotsiri2018]. The irreversible nature of settlement ensures that once a cyprid commits to a location, it cannot relocate, making settlement decisions crucial for long-term survival and reproductive success. This biological constraint provides fundamental insights for designing underwater robotic systems that must balance exploration with permanent deployment strategies in dynamic marine environments where relocation is energetically costly and communication is intermittent [wright1999influence].
@@ IV-B Larval Drift and Hydrodynamic Cues
Barnacle larvae progress through six naupliar stages before reaching the cyprid phase, consistent with observations in both intertidal and epizoic barnacle species. The development time is temperature-dependent, with cyprids emerging in approximately 7 to 8 days at 20 to 25 °C, followed by settlement and metamorphosis reaching juvenile form on day 12 to 18. During this phase of planktonic drift, larvae are passively transported by ambient currents [nogata2006larval], enabling broad dispersal before settlement.
When approaching large marine hosts, hydrodynamics strongly shapes micro-habitats of encounter and settlement. The larvae actively swim to exploit the favourable flow structures for settlement [larsson2016instantaneous]. Recent work by Reidenbach et al. demonstrates hydrodynamic interactions with coral topography and their impact on larval settlement [reidenbach2021hydrodynamic], while Ella and Genin examine capture dynamics at different flow speeds [ella2024capture].
Analyses on small cetaceans show that the commensal barnacle Xenobalanus globicipitis concentrates along the trailing edges of the appendages, with preference for the centre of the fluke. In delphinids, the densities of the dorsal side exceed the densities of the ventral side, consistent with asymmetric fluke kinematics and adhesion-related shear forces [carrillo2015living, Pugliese2012]. The trailing-edge colonisation pattern can be observed directly on the dorsal fins of killer whales, as shown in Fig. 4.
Figure 4: Photographs of four killer whales (O. orca) observed off northern Baffin Island in the eastern Canadian Arctic with Xenobalanus barnacles attached to the trailing edge of their upper dorsal fins. Adapted from Matthews [matthews2020epizoic].
@@ IV-C Host Selection and Settlement
Beyond hydrodynamic guidance, chemical confirmation is a key determinant of settlement. Recent work identified a waterborne settlement pheromone for barnacles as adenosine. Settlement activity remains robust to experimental warming and acidification, underscoring the reliability of chemical signalling under fluctuating conditions [Wu2024].
Recent studies demonstrate that exogenous adenosine significantly increases cyprid settlement rates and exploration frequency. It also improves the mechanical properties of adhesive footprints [xu2025exogenous]. Liu et al. examine chemical signalling in biofilm-mediated bio-fouling [liu2024chemical], providing information on settlement mechanisms.
In practice, cyprids integrate hydrodynamic concentration near suitable micro-habitats with host-derived chemical signals [wright1999influence]. Settlement occurs only when cue strength and contact time exceed a threshold, as anchoring is irreversible.
@@ IV-D Colony Formation and Influence
Following anchoring, local hydrodynamics and conspecific signalling reinforce gregarious settlement. This drives colony growth along flow-favourable edges [mignucci2022barnacles]. Positive feedback improves reproductive opportunity and feeding efficiency through consistent flow delivery. It imposes negligible drag costs for large hosts compared to body mass.
Host specificity and adaptive evolution in settlement behaviour further enhance colony formation, as demonstrated in coral-associated barnacle larvae [yap2023host]. Colonisation patterns such as those shown in Fig. 4 provide the biological basis for NOAH’s anchoring and colony-communication modules.
@@ V-A State and Fields
which combines energy cost, risk exposure, coverage gaps, and information gain into a single score [deb2011multi].
@@
where $\tilde{v}_{i,t+1}$ is the tentative velocity update. The prefactor $(1-a_{i,t})$ ensures settled agents remain frozen, implementing irreversible anchoring analogous to cyprid larvae [kitade2022faint]. This progressive freeze strategy transitions agents from exploration to exploitation [kirkpatrick1983optimization, cao2017optimization], extending multi-swarm PSO frameworks [liang2005dynamic] with irreversible anchoring.
@@ V-B Drift and Update Equations
Anchored colonies emit a spatial field that is repulsive in the short range (to avoid overcrowding) [hooper2016spatial], but attractive in the longer range (to draw agents into the vicinity). This bio-fouling pressure field mimics how barnacle colonies signal both attraction and exclusion zones, similar to how real barnacles use chemical cues to guide larval settlement [Kotsiri2018]. The field incorporates communication noise through probabilistic signal transmission (Equation 4), defined as
@@
where $A\in[0.1,1.0]$ and $B\in[0.1,0.5]$ control the attraction and repulsion strengths, while $\sigma_{a}\in[0.2,0.8]$ and $\sigma_{r}\in[0.05,0.3]$ define the spatial scales (with $\sigma_{a}>\sigma_{r}$). The communication parameters $R_{k}\in[0.1,0.5]$ represent the reliable communication range for the colony $k$, and $\alpha\in[1.0,5.0]$ controls signal attenuation beyond this range, reflecting realistic underwater acoustic propagation characteristics [connor2021current].
The parameter ranges reflect biological observations. Attraction strength $A$ corresponds to chemical signalling intensity in barnacle colonies, while repulsion strength $B$ models physical crowding constraints. The spatial scales $\sigma_{a}$ and $\sigma_{r}$ mirror the effective range of chemical signals versus the zones of immediate physical contact observed in marine larvae [wright1999influence].
Based on cyprid swimming speeds of $\sim$1.8 cm/s [larsson2016instantaneous, Kotsiri2018], the attraction ranges significantly exceed the immediate contact zones. This enables larvae to detect and navigate the appropriate settlement sites from considerable distances. The dual-scale approach ensures that agents are drawn toward productive areas, but avoid overcrowding. This mirrors the biological phenomenon where barnacle larvae are attracted to established colonies but settle at appropriate distances to avoid competition for resources.
The gradient $G_{\Phi}(x)=\nabla\Phi(x)$ acts as a virtual force on free agents, guiding their movement toward optimal settlement regions while maintaining spatial diversity [gazi2004class, gazi2006coordination]. This dual-scale interaction mirrors empirical observations in barnacle larvae, where cyprids are chemically attracted to productive conspecific aggregations, but still reject immediate-contact regions to avoid overcrowding [Berntsson2004]. Furthermore, larvae can detect dissolved settlement signals under flow before making direct contact with the substrate, allowing field-driven navigation rather than purely contact-based adhesion [Koehl2004]. The Gaussian-based potential field approach follows established methods of the artificial potential field for swarm robotics [baziyad2023direction], while the communication probability model incorporates realistic characteristics of the wireless link observed in transition regions [zuniga2004analyzing]. The attraction-repulsion dynamics are based on social force models [helbing1995social] (see Equation 4).
@@
The velocity update mechanism integrates five biologically inspired components that mirror how barnacle nauplii navigate within hydrodynamic environments. Let $g(x_{i,t})$ denote a surrogate estimate of the local gradient of $f$ at the position of agent $i$. This estimate may be obtained through finite differences or relative fitness ranks while keeping NOAH derivative-free. The tentative velocity update $\tilde{v}_{i,t+1}$ is given by:
@@
$\quad+\underbrace{\beta\left(-g(x_{i,t})\right)}_{\text{fitness gradient}}+\underbrace{\gamma U(x_{i,t})}_{\text{drift (flow)}}$
@@
where $\xi_{i,t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ introduces stochastic exploration. The inertia term $(\omega v_{i,t})$ follows the particle swarm optimisation framework [shi1998modified], where the inertia weight $\omega$ balances global and local search capabilities [shi2001comparing]. The random component $(\eta\xi_{i,t})$ employs Gaussian perturbations consistent with stochastic approximation theory, particularly simultaneous perturbation stochastic approximation methods [spall1999adaptive, maryak2001global]. The fitness gradient term $(\beta(-\hat{\nabla}f))$ uses finite difference approximations for derivative-free optimisation [boresta2022mixed, khanh2023general], while the drift term $(\gamma U(x_{i,t}))$ models the hydrodynamic flow effects observed in ocean sampling applications [paley2008cooperative]. Finally, the colony field $(\delta G_{\Phi}(x_{i,t}))$ incorporates the mean-field control principles for swarm coordination [carra2017coalescing, cui2024density].
The parameter ranges are biologically motivated based on observed swimming speeds of 2.6 to 15.2 body lengths per second [wong2020swimming] and rheotactic attachment behaviours in which 93% of cyprids swim upstream before settlement [larsson2016instantaneous]. Momentum retention $\omega\in[0.5,0.9]$ reflects larval inertia, exploration parameter $\eta\in[0.1,0.5]$ models stochastic swimming variability, gradient following strength $\beta\in[0.2,1.0]$ corresponds to chemotactic responsiveness, current adaptation $\gamma\in[0.3,0.8]$ reflects hydrodynamic flow exploitation, and colony influence $\delta\in[0.1,0.6]$ balances attraction to settlement-rich regions with localised spacing effects [carrillo2015living].